# High-dimensional quantum key distribution using a multi-plane light converter

### v1

@@
High-dimensional quantum key distribution (QKD) offers higher information capacity and stronger resilience to noise compared to its binary counterpart. However, these advantages are often hindered by the difficulty of realizing the required high-dimensional measurements and transformations. Here, we implement a large-scale multi-plane light converter (MPLC) and program it as a high-dimensional mode sorter of spatial modes for QKD. We use the MPLC to demonstrate five-dimensional QKD using six mutually unbiased bases and 25-dimensional QKD with two mutually unbiased bases in the same experimental setup. We further propose a construction of pairs of mutually unbiased bases that are robust to experimental errors, and their measurement complexity scales only with the square root of the encoded dimension, paving the way for QKD implementations in higher dimensions.
@@ 1 Introduction
A promising approach for increasing the information capacity of QKD systems and improving their resilience to noise is through high-dimensional encoding[7, 8, 9]. In these systems, high-dimensional degrees of freedom are used to encode d-dimensional quantum states (qudits) that are then measured in two or more mutually unbiased bases (MUBs). High-dimensional quantum key distribution protocols have been demonstrated utilizing spatial [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], time-bin[25, 26, 27, 28, 29, 30], or time-energy [31, 32, 33, 34, 35, 36, 37, 38, 39] encoding. Spatial encoding is particularly attractive for high-dimensional QKD, as spatial light modulators (SLMs) allow efficient control over such high-dimensional states, which can then be transmitted through either multimode fibers[17, 40, 41], multi-core fibers[21, 23], or free-space links[42, 14].
@@
In this work, we use a record-large programmable MPLC with 10 phase planes to sort multiple MUBs of spatially entangled photons for high-dimensional QKD. We demonstrate two high-dimensional QKD schemes using the same experimental setup. We first implement a five-dimensional QKD scheme with sorted measurements in all six MUBs, obtaining a secure key rate of 1.15 bits per sifted photon. We then perform 25-dimensional QKD in two MUBs. We take advantage of the flexibility and programmability of the MPLC to circumvent the fact that arbitrary mode sorters in a 25-dimensional space are beyond state-of-the-art experimental capabilities. Instead of the standard computational and discrete Fourier transform (DFT) bases, we use the MPLC to measure the photons in tailored MUBs with a complexity that scales only with the square root of the dimension. We further show that the structure of such MUBs leads to a non-uniform experimental error distribution, which is beneficial to the secure key rate, measured to be 0.81 bits per sifted photon in the experiment.
@@ 2 Mode sorting with a multi-plane light converter
In our experiment, we use the MPLC for high-dimensional QKD (fig.1a, see Supplementary). Alice produces pairs of spatially entangled photons via spontaneous parametric down-conversion (SPDC)[52]. The SPDC light passes at the far field through a binary amplitude mask with 50 apertures that define a 25-dimensional ’pixel-entangled’ quantum state[53]. To encode the information, Alice chooses a MUB by programming the appropriate unitary transformation between the set of input and output spatial modes using the MPLC. She then measures her photon at one of the output modes, which determines the state of the photon sent to Bob. To decode the information, Bob uses a second MPLC to sort the modes in a chosen MUB and measures his photon at the output. For experimental convenience, we realize the two independent MPLCs by bouncing the entangled photons ten times between a single SLM and a mirror, where the top part of each phase mask is used by Alice and the bottom part by Bob.
Figure 1: Spatially-encoded QKD using multi-plane light conversion. (a) Pairs of spatially entangled photons are generated via SPDC and are transmitted through a binary amplitude mask with 50 apertures. Alice uses a programmable MPLC with ten phase masks to measure one of the photons in one out of several MUBs. Its twin photon is transmitted to Bob, who uses his programmable MPLC to sort the different spatial modes in one of the MUBs. Measurements in which Alice and Bob choose the same basis contribute to the establishment of a secret key. Examples of such measurements in two MUBs of dimension 25 are presented in panels (b) and (c), where the optimized MPLC phase masks and the sum of intensities from all input modes are plotted for Alice (top half) and Bob (bottom half) for all planes.
@@ 3 Quantum key distribution in six mutually unbiased bases
To demonstrate the flexibility of MPLCs, we first consider five spatial modes per photon and perform five-dimensional QKD with measurements in all six MUBs[7, 54]. Along with the standard computational basis {|$\{\ket{n}\}_{n=1,...,5}$ that can be measured without the MPLC, we choose five other MUBs {|$\left\{\ket{\psi^{(r)}_{k}}\right\}_{k=1,...,5}$, with $k$ indexing the state and $r=2,...,6$ the basis, such that[55]
@@
For $r=2$, Eq. 1 is exactly the Discrete Fourier Transform (DFT) basis of dimension five. Note that the four other bases $r=3,..,6$ only differ from the DFT basis by phases that are independent of the state $k$. Therefore, bases $r=3,..,6$ can be mapped onto the DFT basis by applying mode-dependent phases to the spatial modes. For our pixel-entangled states, this is easily implemented using a single phase mask at the first plane of the MPLC. Measurements in these MUBs are thus realized by encoding the DFT transformation using the MPLC and superimposing mode-dependent phases at the first plane of the MPLC to choose the desired MUB (fig.2a).
The probabilities for each of Bob’s measurement outcomes given a state sent by Alice, in all MUBs, are presented in fig.2b. We obtain an average error rate of $E=11\%$. The secure key rate is lower bounded by $R=log_{2}(d)-h_{d}(\frac{d+1}{d}E)-\frac{d+1}{d}Elog_{2}(d+1)=1.15$ bits per sifted photon, where $h_{d}(x)=-xlog_{2}(\frac{x}{d-1})-(1-x)log_{2}(1-x)$ is the high-dimensional Shannon entropy[7, 56, 16].
Figure 2: Five-dimensional QKD using a mode sorter in six MUBs. We program the MPLC to sort the photons in all six MUBs of dimension five. The bases are chosen such that one basis is the trivial computational basis, while the others are measured by adding different relative phases at the first plane of the MPLC, before realizing an optical discrete Fourier transform (DFT)(a). We measure the correlations between the photons in all bases and calculate the probability of each outcome in Bob’s measurement for each state sent by Alice (b). From these measurements, we extract a secure key rate of 1.15 bits per sifted photon.
@@ 4 Experimentally-efficient 25-dimensional quantum key distribution
Figure 3: 25-dimensional QKD in two MUBs. We encode a 25-dimensional quantum state in a grid of 25 spatially non-overlapping modes. Instead of using the standard computational and 25-dimensional DFT bases, we design two MUBs that are tailored to the experimental capabilities and only require the interference of sets of five spatial modes (a),(b). Measurements in these MUBs are performed by applying five-dimensional DFT transformations along the rows (a) or columns (b) of the 5x5 grid of modes using the MPLC. The experimental results are presented in (c), yielding a secure key rate of 0.81 bits per photon.
Interestingly, a non-uniform error distribution is clearly visible from the correlations in fig.3c. Along with uniform error sources such as accidental coincidence counts, the structure of the unitary transformation, that interferes blocks of $\sqrt{d}$ modes, leads to a block-biased error distribution. Non-uniform error distributions are also common in other mode sets such as Lagurre Gaussian modes in free space [59], linearly polarized modes in multimode fibers [60] or even time-bins [30]. In all cases, the block-biased noise behavior stems from the mode and MUB structures that dictate a lower probability of errors for distant modes that do not interfere. In what follows, we utilize this structure to get a tighter lower bound for the secure key rate.
@@
$\begin{split}h_{d}(E_{u},E_{b})=-(1-E_{t})log_{2}(1-E_{t})-\left(d-\sqrt{d}\right)\frac{E_{u}}{d-1}log_{2}\left(\frac{E_{u}}{d-1}\right)\\ -\left(\sqrt{d}-1\right)\left(\frac{E_{b}}{\sqrt{d}-1}+\frac{E_{u}}{d-1}\right)log_{2}\left(\frac{E_{b}}{\sqrt{d}-1}+\frac{E_{u}}{d-1}\right)\end{split}$ (2)
In Eq. 2, the first term is for the measurement of the correct state, the second is for the $d-\sqrt{d}$ states outside the block, and the third is for the $\sqrt{d}-1$ possible error states within the block that are affected by both the uniform and block errors. Since uniform error maximizes the Shannon entropy, any non-uniformity in the error distribution increases the achievable secure key rate and total tolerable noise (fig.4a,b). The secure key rate is lower-bounded by $R=log_{2}(d)-2h_{d}(E_{u},E_{b})$[61, 56], yielding 0.81 secure bits per sifted photon in our experiment.
Figure 4: Secret key rates and error threshold. (a) The secret key rate per sifted photon in QKD is dependent on both the total error rate as well as the error distribution. In our 25-dimensional protocol, as a direct result of the block-like structure of the optical transformations, the errors are clearly not uniform and exhibit a strong block-like bias (orange curve). We therefore use a block-biased error model to calculate the secure key rate in this case, obtaining a rate of 0.81 bits per sifted photon (black marker). Non-uniform errors also increase the error threshold of the protocol (b). The error threshold increases significantly when the errors are block-biased, as in our experiment (orange), and even more so when they are completely within the block (green).
@@ 5 Conclusion
For dimensions that are an even power of a prime number, we have further proposed pairs of MUBs for which the complexity of the experimental transformations only scales with the square root of the dimension. We have performed measurements in such MUBs for dimension 25, and shown that they lead to a non-uniform error distribution that benefits the secure key rate.
@@
We thus believe that incorporating our demonstration of high-dimensional QKD using MPLCs into state-of-the-art spatial high-dimensional QKD systems [10, 11, 12, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] would pave the way towards high-rate long-range QKD.
@@
Funding This research was funded by the Israeli Innovation Authority (Quantum Communication consortium) and the Israel Science Foundation (grant No. 2497/21). O.L. acknowledges the support of the Clore Scholars Programme of the Clore Israel Foundation. K.S. acknowledges the support of the Israeli Council for Higher Education. Y.B. and K.S. acknowledge the support of the Zuckerman STEM Leadership Program.

### v2

@@
High-dimensional quantum key distribution (QKD) offers higher information capacity and stronger resilience to noise compared to its binary counterpart. However, these advantages are often hindered by the difficulty of realizing the required high-dimensional measurements and transformations. Here, we implement a large-scale multi-plane light converter (MPLC) and program it as a high-dimensional mode sorter of spatial modes for QKD. Using the MPLC, we demonstrate five-dimensional QKD with six mutually unbiased bases and 25-dimensional QKD with two mutually unbiased bases in the same experimental setup. Furthermore, we propose a construction of pairs of mutually unbiased bases that are robust to experimental errors, with measurement complexity scaling only with the square root of the encoded dimension. This approach paves the way for QKD implementations in higher dimensions.
@@ 1 Introduction
A promising approach for increasing the information capacity of QKD systems and improving their resilience to noise is through high-dimensional encoding[7, 8, 9]. In these systems, high-dimensional degrees of freedom are used to encode d-dimensional quantum states (qudits) that are then measured in two or more mutually unbiased bases (MUBs). High-dimensional quantum key distribution protocols have been demonstrated utilizing spatial [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], time-bin[26, 27, 28, 29, 30, 31], or time-energy [32, 33, 34, 35, 36, 37, 38, 39, 40] encoding. Spatial encoding is particularly attractive for high-dimensional QKD, as spatial light modulators (SLMs) allow efficient control over such high-dimensional states, which can then be transmitted through either multimode fibers[17, 41, 42], multi-core fibers[21, 23], or free-space links[43, 14].
@@
In this work, we use a record-large programmable MPLC with 10 phase planes to sort multiple MUBs of spatially entangled photons for high-dimensional QKD. We demonstrate two high-dimensional QKD schemes using the same experimental setup. We first implement a five-dimensional QKD scheme with sorted measurements in all six MUBs, obtaining a secure key rate of 1.57 bits per sifted photon through semidefinite programming on the tomographically-complete data[53]. We then perform 25-dimensional QKD in two MUBs. We take advantage of the flexibility and programmability of the MPLC to circumvent the fact that arbitrary mode sorters in a 25-dimensional space are beyond state-of-the-art experimental capabilities. Instead of the standard computational and discrete Fourier transform (DFT) bases, we use the MPLC to measure the photons in tailored MUBs with a complexity that scales only with the square root of the dimension. We further show that the structure of such MUBs leads to a non-uniform experimental error distribution, which is beneficial to the secure key rate, measured to be 0.8 bits per sifted photon in the experiment.
@@ 2 Mode sorting with a multi-plane light converter
In our experiment, we use the MPLC for high-dimensional QKD (fig.1a, see Supplementary). Alice produces pairs of spatially entangled photons via spontaneous parametric down-conversion (SPDC)[54]. The SPDC light passes at the far field through a binary amplitude mask with 50 apertures that define a 25-dimensional ’pixel-entangled’ quantum state[55]. To encode the information, Alice chooses a MUB by programming the appropriate unitary transformation between the set of input and output spatial modes using the MPLC. She then measures her photon at one of the output modes, which determines the state of the photon sent to Bob. To decode the information, Bob uses a second MPLC to sort the modes in a chosen (transposed) MUB and measures his photon at the output. For experimental convenience, we realize the two independent MPLCs by bouncing the entangled photons ten times between a single SLM and a mirror, where the top part of each phase mask is used by Alice and the bottom part by Bob.
Figure 1: Spatially-encoded QKD using multi-plane light conversion. (a) Pairs of spatially entangled photons are generated via SPDC and are transmitted through a binary amplitude mask with 50 apertures. Alice uses a programmable MPLC with ten phase masks to measure one of the photons in one of several MUBs. Its twin photon is transmitted to Bob, who uses his programmable MPLC to sort the different spatial modes in one of the MUBs. Measurements in which Alice and Bob choose the same basis contribute to the establishment of a secret key. Examples of such measurements in two MUBs of dimension 25 are presented in panels (b) and (c), where the optimized MPLC phase masks and the sum of intensities from all input modes are plotted for Alice (top half) and Bob (bottom half) for all planes.
@@ 3 Quantum key distribution in six mutually unbiased bases
To demonstrate the flexibility of MPLCs, we first consider five spatial modes per photon and perform five-dimensional QKD with measurements in all six MUBs[7, 56]. Along with the standard computational basis {|$\{\ket{n}\}_{n=1,\ldots,5}$ that can be measured without the MPLC, we choose five other MUBs {|$\left\{\ket{\psi^{(r)}_{k}}\right\}_{k=1,\ldots,5}$, with $k$ indexing the state and $r=2,\ldots,6$ the basis, such that[57]
@@
For $r=2$, Eq. 1 is exactly the Discrete Fourier Transform (DFT) basis of dimension five. Note that the four other bases $r=3,\ldots,6$ only differ from the DFT basis by phases that are independent of the state $k$. Therefore, bases $r=3,\ldots,6$ can be mapped onto the DFT basis by applying mode-dependent phases to the spatial modes. For our pixel-entangled states, this is easily implemented using a single phase mask at the first plane of the MPLC. Measurements in these MUBs are thus realized by encoding the DFT transformation using the MPLC and superimposing mode-dependent phases at the first plane of the MPLC to choose the desired MUB (fig.2a).
The estimated probabilities for each of Bob’s measurement outcomes given a state sent by Alice, in all MUBs, are presented in fig.2b. From these probabilities, it is possible to compute a lower bound on the secure key rate. A rough lower bound can be obtained by applying the often-used simplifying assumption of identical depolarizing noise channels in all bases[7, 58, 16, 59]. This yields a secure key rate of $R=\log_{2}(d)-h_{d}(\frac{d+1}{d}E)-\frac{d+1}{d}E\log_{2}(d+1)=1.15\pm 0.05$ bits per sifted photon, where $h_{d}(x)=-x\log_{2}(\frac{x}{d-1})-(1-x)\log_{2}(1-x)$ is the high-dimensional Shannon entropy and $E=11\%$ is the arithmetic mean of the error rates in all bases.
However, a tighter estimation of the lower bound could be obtained through numerical techniques[60, 61, 31, 62, 53] using the full tomographically-complete data in fig.2b[53]. Indeed, using the method described in [53], we observe a gradual increase in the estimated key rate when taking into account more of the experimental data, both when Alice and Bob measure in the same basis and when measuring in different bases (see Supplementary information). We obtain a secure key rate of 1.57 for the full tomographically-complete data, which is significantly higher than the loose bound given by assuming depolarizing noise.
Figure 2: Five-dimensional QKD using a mode sorter in six MUBs. We program the MPLC to sort the photons in all six MUBs of dimension five. The bases are chosen such that one basis is the trivial computational basis, while the others are measured by adding different relative phases at the first plane of the MPLC, before realizing an optical discrete Fourier transform (DFT)(a). We measure the correlations between the photons in all bases and estimate the probability of each outcome in Bob’s measurement for each state sent by Alice (b). From these tomographically-complete measurements, we extract a secure key rate of 1.57 bits per sifted photon.
@@ 4 Experimentally-efficient 25-dimensional quantum key distribution
Figure 3: 25-dimensional QKD in two MUBs. We encode a 25-dimensional quantum state in a grid of 25 spatially non-overlapping modes. Instead of using the standard computational and 25-dimensional DFT bases, we design two MUBs that are tailored to the experimental capabilities and only require the interference of sets of five spatial modes (a),(b). Measurements in these MUBs are performed by applying five-dimensional DFT transformations along the rows (a) or columns (b) of the 5x5 grid of modes using the MPLC. The experimental results are presented in (c), yielding a secure key rate of 0.8 bits per photon.
Interestingly, a non-uniform error distribution is clearly visible from the correlations in fig.3c. Along with uniform error sources such as accidental coincidence counts, the structure of the unitary transformation, that interferes blocks of $\sqrt{d}$ modes, leads to a block-biased error distribution. Non-uniform error distributions are also common in other mode sets such as Lagurre Gaussian modes in free space [65], linearly polarized modes in multimode fibers [66] or even time-bins [31]. In all cases, the block-biased noise behavior stems from the mode and MUB structures that dictate a lower probability of errors for distant modes that do not interfere. As the complexity of the semidefinite programming approach we used for $d=5$ prevents its implementation for $d=25$, in what follows, we utilize this block-biased noise structure to get a tighter analytical lower bound for the secure key rate.
@@
$\begin{split}h_{d}(E_{u},E_{b})=-(1-E_{t})\log_{2}(1-E_{t})-\left(d-\sqrt{d}\right)\frac{E_{u}}{d-1}\log_{2}\left(\frac{E_{u}}{d-1}\right)\\ -\left(\sqrt{d}-1\right)\left(\frac{E_{b}}{\sqrt{d}-1}+\frac{E_{u}}{d-1}\right)\log_{2}\left(\frac{E_{b}}{\sqrt{d}-1}+\frac{E_{u}}{d-1}\right)\end{split}$ (2)
In Eq. 2, the first term is for the measurement of the correct state, the second is for the $d-\sqrt{d}$ states outside the block, and the third is for the $\sqrt{d}-1$ possible error states within the block that are affected by both the uniform and block errors. Since uniform error maximizes the Shannon entropy, any non-uniformity in the error distribution increases the achievable secure key rate and total tolerable noise (fig.4a,b). The secure key rate is lower-bounded by $R=\log_{2}(d)-2h_{d}(E_{u},E_{b})$[67, 58, 59], yielding $0.8\pm 0.07$ secure bits per sifted photon in our experiment.
Figure 4: Secret key rates and error threshold. (a) The secret key rate per sifted photon in QKD is dependent on both the total error rate as well as the error distribution. In our 25-dimensional protocol, as a direct result of the block-like structure of the optical transformations, the errors are clearly not uniform and exhibit a strong block-like bias (orange curve). We therefore use a block-biased error model to calculate the secure key rate in this case, obtaining a rate of 0.8 bits per sifted photon (black marker). Non-uniform errors also increase the error threshold of the protocol (b). The error threshold increases significantly when the errors are block-biased, as in our experiment (orange), and even more so when they are completely within the block (green).
@@ 5 Conclusion
For dimensions that are an even power of a prime number, we have further proposed pairs of MUBs for which the complexity of the experimental transformations only scales with the square root of the dimension. We have performed measurements in such MUBs for dimension 25 and shown that they lead to a non-uniform error distribution that allows for obtaining a tighter analytical lower bound for the secure key rate.
@@
We thus believe that incorporating our demonstration of high-dimensional QKD using MPLCs into state-of-the-art spatial high-dimensional QKD systems [10, 11, 12, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 69, 24, 70] would pave the way towards high-rate long-range QKD.
@@
Funding This research was funded by the Israeli Innovation Authority (Quantum Communication Consortium) and the Israel Science Foundation (grants No. 2497/21 and 2137/19). O.L. acknowledges the support of the Clore Scholars Programme of the Clore Israel Foundation. K.S. acknowledges the support of the Israeli Council for Higher Education. Y.B. and K.S. acknowledge the support of the Zuckerman STEM Leadership Program. The research of M.A. was supported by the European Union–Next Generation UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León.